Solve the IVP. 1 [- [ 3 ] [:]+[66*) ()-14
Answer #1

[31]11-[*] [101-1 so, x=-5x+3y + 6tözt -2t 18té 1 +9 -3 (-50+3y + 6 tezt) + y = -5(-36) - gy- 18 tetty -5(y-7) - gy-1Bt -5824 제= ty 는 -24 ++ (-2zit + -( -22 분 + tttg t란 3러는 = -29호, 2월4호랑도장, 랑 활량 9많라 2달란 3관한 Q 라 2 -925 tte3란라, 9을라 기 올랑 란 용호란 , 3종 =